A posion distribution has double mode at x=2 and x=3 .what is the probability that x will have one or other of the two values
1
Expert's answer
2021-10-10T18:26:54-0400

If the Poisson distribution is bimodal, the two modes are at the points x=λ-1 and x=λ, where λ is the parameter of the Poisson distribution.

"x=1 \\

x=2 \\

λ=2 \\

P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

P(X=1) = 2e^{-2} \\

P(X=2) = \frac{e^{-2}2^2}{2!} = 2e^{-2}"

Required probability "= P(X=1) + P(X=2) = 2e^{-2} +2e^{-2} = 0.542"